Please solve by factors
4x squared -29x+7=0



Thank you. I had the answer but lost it.

Answer by stanbon(75887)

solve by factors
4x^2 -29x+7 = 0
-------
4x^2 - 28x - x + 7 = 0
---
4x(x-7)-(x-7) = 0
(x-7)(4x-1) = 0
x = 7 or x = 1/4
=================
Cheers,
Stan H.